One of two supplementary angles measures 37 degrees less than four times the other. Find the measure of each angle.
Answer by bam878s(77)

we know supplementary angles add up to 180 degrees
so for angle 1 (a1) and angle 2 (a2)
we have a1+a2=180 and a1=4a2-37
sub a1 in the first equation to get
4a2-37+a2=180
add 37 to both sides to get
5a2=217
divide both sides bt 5
a2=43.4
now sub this back in the first equation
a1=180-a2=180-43.4=136.6